OpenStudy (credmond):

Using standard potential, calculate the standard cell potential and the equilibrium constants for the following reactions: Cu(s) + 2Ag+(aq) -> Cu2+(aq) + 2Ag(s) Zn(s) + Fe2+(aq) -> Zn2+(aq) + Fe(s)

OpenStudy (rina.r):

i think that's how u do for A a. Cu²⁺(aq) + 2e⁻ ⇌ Cu(s) ; E° = 0.340V Ag⁺(aq) + e⁻ ⇌ Ag(s) ; E° = 0.7996V => Cu(s) ⇌ Cu²⁺(aq) + 2 e⁻ ; E° = -0.34V 2Ag⁺(aq) + 2e⁻ ⇌ Ag(s) ; E° = 2∙7996V = 1.5992V => Cu(s) + 2 Ag⁺(aq) ⇌ Cu²⁺(aq) + 2 Ag(s) ; E° = -0.340V - 1.5992.4V = 1.2592V => K = e^(38.941V⁻¹ ∙ 2 ∙ 1.2592V ) = e^( 98.069) = 3.90×10⁴²